OpenStudy (anonymous):

is it possible to create a fourth-degree polynomial with only two real zeros. Demonstrate how to do this and explain your steps.

OpenStudy (anonymous):

It is possible, if two zeros are real numbers and the other two zeros are complex numbers. Example: (x-1)(x+1)(x^2 + 1) = 0 or (x^2-1)(x^2+1) = 0 The real zeros are 1 and -1; the complex zeros are i and - 1. Because when you set x^2 + 1 = 0, you get x^2 = -1, and only i and -i are solutions. x^2 - 1 = 0 has two real zeros, x = 1 and -1.

OpenStudy (anonymous):

thank you